One angle of a triangle measures 6 degrees more than the second. The third angle measures 4 degrees less than 3 times of the sum of the measures of the first two angles. Find the measures of the three angles.
The first angle measures____ degrees, the second_____ degrees, and the third_____degrees.

Answer by checkley71(8403)

let x be the first angle
let x+6 be the next angle
let 3(x+x+6)
x+x+6+3(x+x+6)=180
2x+6+3(2x+6)=180
2x+6+6x+18=180
8x=180-6-18
8x=156
x=156/8
x=19.5 degrees for the smallest angle.
19.5+6=25.5 degrees for the second angle.
3(19.5+25.5)
3(45)=145 the largest angle.
proof
19.5+25.5+145=180
180=180